# One-Step Effective Diffusion Network for Real-World Image Super-Resolution

### v1

@@ Abstract
The pre-trained text-to-image diffusion models have been increasingly employed to tackle the real-world image super-resolution (Real-ISR) problem due to their powerful generative image priors. Most of the existing methods start from random noise to reconstruct the high-quality (HQ) image under the guidance of the given low-quality (LQ) image. While promising results have been achieved, such Real-ISR methods require multiple diffusion steps to reproduce the HQ image, increasing the computational cost. Meanwhile, the random noise introduces uncertainty in the output, which is unfriendly to image restoration tasks. To address these issues, we propose a one-step effective diffusion network, namely OSEDiff, for the Real-ISR problem. We argue that the LQ image contains rich information to restore its HQ counterpart, and hence the given LQ image can be directly taken as the starting point for diffusion, eliminating the uncertainty introduced by random noise sampling. We finetune the pre-trained diffusion network with trainable layers to adapt it to complex image degradations. To ensure that the one-step diffusion model could yield HQ Real-ISR output, we apply variational score distillation in the latent space to conduct KL-divergence regularization. As a result, our OSEDiff model can efficiently and effectively generate HQ images in just one diffusion step. Our experiments demonstrate that OSEDiff achieves comparable or even better Real-ISR results, in terms of both objective metrics and subjective evaluations, than previous diffusion model based Real-ISR methods that require dozens or hundreds of steps. The source codes will be released at https://github.com/cswry/OSEDiff.
@@ 1 Introduction
Image super-resolution (ISR) [12, 64, 28, 63, 6, 23, 45, 26, 59] is a classical yet still active research problem, which aims to restore a high-quality (HQ) image from its low-quality (LQ) observation suffering from degradations of noise, blur and low-resolution, etc. While one line of ISR research [12, 64, 28, 63, 6] simplifies the degradation process from HQ to LQ images as bicubic downsampling (or downsampling after Gaussian blur) and focus on the study on network architecture design, the trained models can hardly be generalized to real-world LQ images, whose degradations are often unknown and much more complex. Therefore, another increasingly popular line of ISR research is the so-called real-world ISR (Real-ISR) [59, 44] problem, which aims to reproduce perceptually HQ images from the LQ images collected in real-world applications.
@@
Figure 1: Performance and efficiency comparison among SD-based Real-ISR methods. (a). Performance comparison on the DrealSR benchmark [50]. Metrics like LPIPS and NIQE, where smaller scores indicate better image quality, are inverted and normalized for display. OSEDiff achieves leading scores on most metrics with only one diffusion step. (b). Model efficiency comparison. The inference time is tested on an A100 GPU with $512\times 512$ input image size. OSEDiff has the fewest trainable parameters and is over 10 times faster than the recently developed SeeSR [51].
@@
To address the aforementioned issues, we propose a One-Step Effective Diffusion network, OSEDiff in short, for the Real-ISR problem. The UNet backbone in pre-trained SD models has strong capability to transfer the input data into another domain, while the given LQ image actually has rich information to restore its HQ counterpart. Therefore, we propose to directly feed the LQ images into the pre-trained SD model without introducing any random noise. Meanwhile, we integrate trainable LoRA layers [16] into the pre-trained UNet, and finetune the SD model to adapt it to the Real-ISR task. On the other hand, to ensure that the one-step model can still produce HQ natural images as the multi-step models, we utilize variational score distillation (VSD) [48]
for KL-divergence regularization. This operation effectively leverages the powerful natural image priors of pre-trained SD models and aligns the distribution of generated images with natural image priors. As illustrated in Fig. 1, our extensive experiments demonstrate that OSEDiff achieves comparable or superior performance measures to state-of-the-art SD-based Real-ISR models, while it significantly reduces the number of inference steps from N to 1 and has the fewest trainable parameters, leading to more than ×$\times 10$ speedup in inference time than the recent state-of-the-art SeeSR [51].
@@ 3.1 Problem Modeling
However, the performance of such optimization-based methods is largely hindered by two factors. First, the degradation function $\Phi$ is often unknown and hard to model in real-world scenarios. Second, the hand-crafted regularization terms $\mathcal{L}_{\mathrm{reg}}$ are hard to effectively model the complex natural image priors.
With the development of deep-learning techniques, it has become prevalent to learn a neural network $\emph{G}_{\theta}$, which is parameterized by $\theta$, from a training dataset $S$ of $(\boldsymbol{x}_{L},\boldsymbol{x}_{H})$ pairs to map the LQ image to an HQ image. The network training can be described as the following learning problem:
@@
The recently developed T2I diffusion models such as SD [35] offer new options for us to formulate the loss $\mathcal{L}_{\mathrm{reg}}$. These models, trained on billions of image-text pairs, can effectively depict the natural image distribution in latent space. Some score distillation methods have been reported to employ SD to optimize images by using the KL-divergence as the objective [48, 24, 42]. In particular, variational score distillation (VSD) [48] induces such a KL-divergence based objective from particle-based variational optimization to align the distributions represented by two diffusion models. Based on the above discussions, we propose to instantiate the learning objective in Eq. (1) by designing an efficient and effective one-step diffusion network. In specific, we finetune the pre-trained SD with LoRA [16] as our Real-ISR backbone network and employ VSD as our regularizer to align the distribution of network outputs with natural HQ images. The details are provided in the next section.
@@ 3.2 One-Step Effective Diffusion Network
VSD in Latent Space. The original VSD computes the gradients in the image space. When it is used to train an SD-based generator network, there will be repetitive latent decoding/encoding in computing $\mathcal{L}_{\mathrm{reg}}$. This is costly and makes the regularization less effective. Considering the fact that a well-trained VAE should satisfy $\emph{E}_{\phi}(\boldsymbol{x})=\emph{E}_{\phi}(\emph{D}_{\phi}(\boldsymbol{z}))\approx\boldsymbol{z}$, we can approximately let $\emph{E}_{\phi}(\hat{\boldsymbol{x}}_{H})=\hat{\boldsymbol{z}}_{H}$. In this case, we can eliminate the redundant latent encoding/decoding in computing the regularization loss. The gradient of the regularization loss w.r.t. the network parameter $\theta$ in the latent space is:
@@ 4.1 Experimental Settings
Training and Testing Datasets. We train OSEDiff on LSDIR [25] and the first 10K face images of FFHQ [18]. The degradation pipeline of Real-ESRGAN [44] is used to synthesize LQ-HQ training pairs. We evaluate OSEDiff and compare it with competing methods using the test set provided by StableSR [41], including both synthetic and real-world data. The synthetic data includes 3000 images of size $512\times 512$, whose GT are randomly cropped from DIV2K-Val [2] and degraded using the Real-ESRGAN pipeline [44]. The real-world data include LQ-HQ pairs from RealSR [3] and DRealSR [50], with sizes of $128\times 128$ and $512\times 512$, respectively.
@@
Inference Time (s) 11.50 2.72 4.30 2.80 0.71 0.16 0.35
@@
Evaluation Metrics. To provide a comprehensive and holistic assessment on the performance of different methods, we employ a range of full-reference and no-reference metrics. PSNR and SSIM [49] (calculated on the Y channel in YCbCr space) are reference-based fidelity measures, while LPIPS [62], DISTS [11] are reference-based perceptual quality measures. FID [14] evaluates the distance of distributions between GT and restored images. NIQE [60], MANIQA [54], MUSIQ [21], and CLIPIQA [40] are no-reference image quality measures. We also conduct a user study, which is presented in the Appendix.
@@ 4.2 Comparison with State-of-the-Arts
Complexity Comparisons. We further compare the complexity of competing DM-based Real-ISR models in Table 2, including the number of inference steps, inference time, and trainable parameters. All methods are tested on an A100 GPU with an input image of size $512\times 512$. OSEDiff has the fewest trainable parameters, and the trained LoRA layers can be merged into the original SD to further reduce the computational cost. By using only one forward pass, OSEDiff has significant advantage in inference time over multi-step methods. Specifically, OSEDiff is about 28 times faster than StableSR, 10 times faster than SeeSR, and 2 times faster than ResShift. OSEDiff is only slower than another one-step method SinSR, which does not employ the pre-trained SD model. However, the quality of SinSR’s Real-ISR outputs lags significantly behind OSEDiff.
@@ A.4 Algorithm of OSEDiff
The pseudo-code of our OSEDiff training algorithm is summarized as Algorithm 1.
@@ 14      $\hat{\boldsymbol{x}}_{H}\leftarrow\emph{D}_{\theta}(\hat{\boldsymbol{z}}_{H})$
16      $\mathcal{L}_{\mathrm{data}}=\mathcal{L}_{\mathrm{MSE}}(\hat{\boldsymbol{x}}_{H},\boldsymbol{x}_{H})+\lambda_{1}\mathcal{L}_{\mathrm{LPIPS}}(\hat{\boldsymbol{x}}_{H},\boldsymbol{x}_{H})$
      /* Compute regularization objective */
@@ 20      $\hat{\boldsymbol{z}}_{t}\leftarrow\alpha_{t}\hat{\boldsymbol{z}}_{H}+\sigma_{t}\epsilon$
21      $\mathcal{L}_{\mathrm{reg}}\leftarrow\left[\omega(t)(\boldsymbol{\epsilon}_{\theta}(\hat{\boldsymbol{z}}_{t};t,c_{y})-\boldsymbol{\epsilon}_{\theta^{\prime}}(\hat{\boldsymbol{z}}_{t};t,c_{y}))\right]$
      /* Compute regularizer funetuning objective */
22
23      $\mathcal{L}_{\mathrm{diff}}\leftarrow\mathcal{L}_{\mathrm{MSE}}(\boldsymbol{\epsilon}_{\theta^{\prime}}(\hat{\boldsymbol{z}}_{t};t,c_{y}),\epsilon)$

### v3

@@ Abstract
The pre-trained text-to-image diffusion models have been increasingly employed to tackle the real-world image super-resolution (Real-ISR) problem due to their powerful generative image priors. Most of the existing methods start from random noise to reconstruct the high-quality (HQ) image under the guidance of the given low-quality (LQ) image. While promising results have been achieved, such Real-ISR methods require multiple diffusion steps to reproduce the HQ image, increasing the computational cost. Meanwhile, the random noise introduces uncertainty in the output, which is unfriendly to image restoration tasks. To address these issues, we propose a one-step effective diffusion network, namely OSEDiff, for the Real-ISR problem. We argue that the LQ image contains rich information to restore its HQ counterpart, and hence the given LQ image can be directly taken as the starting point for diffusion, eliminating the uncertainty introduced by random noise sampling. We finetune the pre-trained diffusion network with trainable layers to adapt it to complex image degradations. To ensure that the one-step diffusion model could yield HQ Real-ISR output, we apply variational score distillation in the latent space to conduct KL-divergence regularization. As a result, our OSEDiff model can efficiently and effectively generate HQ images in just one diffusion step. Our experiments demonstrate that OSEDiff achieves comparable or even better Real-ISR results, in terms of both objective metrics and subjective evaluations, than previous diffusion model-based Real-ISR methods that require dozens or hundreds of steps. The source codes are released at https://github.com/cswry/OSEDiff.
@@ 1 Introduction
Image super-resolution (ISR) [13, 66, 29, 65, 6, 24, 46, 27, 61] is a classical yet still active research problem, which aims to restore a high-quality (HQ) image from its low-quality (LQ) observation suffering from degradations of noise, blur and low-resolution, etc. While one line of ISR research [13, 66, 29, 65, 6] simplifies the degradation process from HQ to LQ images as bicubic downsampling (or downsampling after Gaussian blur) and focus on the study on network architecture design, the trained models can hardly be generalized to real-world LQ images, whose degradations are often unknown and much more complex. Therefore, another increasingly popular line of ISR research is the so-called real-world ISR (Real-ISR) [61, 45] problem, which aims to reproduce perceptually realistic HQ images from the LQ images captured in real-world applications.
@@
Figure 1: Performance and efficiency comparison among SD-based Real-ISR methods. (a). Performance comparison on the DrealSR benchmark [51]. Metrics like LPIPS and NIQE, where smaller scores indicate better image quality, are inverted and normalized for display. OSEDiff achieves leading scores on most metrics with only one diffusion step. (b). Model efficiency comparison. The inference time is tested on an A100 GPU with $512\times 512$ input image size. OSEDiff has the fewest trainable parameters and is over 100 times faster than StableSR [42].
@@
To address the aforementioned issues, we propose a One-Step Effective Diffusion network, OSEDiff in short, for the Real-ISR problem. The UNet backbone in pre-trained SD models has strong capability to transfer the input data into another domain, while the given LQ image actually has rich information to restore its HQ counterpart. Therefore, we propose to directly feed the LQ images into the pre-trained SD model without introducing any random noise. Meanwhile, we integrate trainable LoRA layers [17] into the pre-trained UNet, and finetune the SD model to adapt it to the Real-ISR task. On the other hand, to ensure that the one-step model can still produce HQ natural images as the multi-step models, we utilize variational score distillation (VSD) [49, 58, 10] for KL-divergence regularization. This operation effectively leverages the powerful natural image priors of pre-trained SD models and aligns the distribution of generated images with natural image priors. As illustrated in Fig. 1, our extensive experiments demonstrate that OSEDiff achieves comparable or superior performance measures to state-of-the-art SD-based Real-ISR models, while it significantly reduces the number of inference steps from N to 1 and has the fewest trainable parameters, leading to more than ×$\times 100$ speedup in inference time over previous methods such as StableSR [42].
@@ 3.1 Problem Modeling
However, the performance of such optimization-based methods is largely hindered by two factors. First, the degradation function $\Phi$ is often unknown and hard to model in real-world scenarios. Second, the hand-crafted regularization terms $\mathcal{L}_{\mathrm{reg}}$ are hard to effectively model the complex natural image priors. With the development of deep-learning techniques, it has become prevalent to learn a neural network $\emph{G}_{\theta}$, which is parameterized by $\theta$, from a training dataset $S$ of $(\boldsymbol{x}_{L},\boldsymbol{x}_{H})$ pairs to map the LQ image to an HQ image. The network training can be described as the following learning problem:
@@
The recently developed T2I diffusion models such as SD [36] offer new options for us to formulate the loss $\mathcal{L}_{\mathrm{reg}}$. These models, trained on billions of image-text pairs, can effectively depict the natural image distribution in latent space. Some score distillation methods have been reported to employ SD to optimize images by using the KL-divergence as the objective [49, 25, 43]. In particular, variational score distillation (VSD) [49, 58, 10] induces such a KL-divergence based objective from particle-based variational optimization to align the distributions represented by two diffusion models. Based on the above discussions, we propose to instantiate the learning objective in Eq. (1) by designing an efficient and effective one-step diffusion network. In specific, we finetune the pre-trained SD with LoRA [17] as our Real-ISR backbone network and employ VSD as our regularizer to align the distribution of network outputs with natural HQ images. The details are provided in the next section.
@@ 3.2 One-Step Effective Diffusion Network
VSD in Latent Space. The original VSD computes the gradients in the image space. When it is used to train an SD-based generator network, there will be repetitive latent decoding/encoding in computing $\mathcal{L}_{\mathrm{reg}}$. This is costly and makes the regularization less effective. Considering the fact that a well-trained VAE should satisfy $\emph{E}_{\phi}(\boldsymbol{x})=\emph{E}_{\phi}(\emph{D}_{\phi}(\boldsymbol{z}))\approx\boldsymbol{z}$, we can approximately let $\emph{E}_{\phi}(\hat{\boldsymbol{x}}_{H})=\hat{\boldsymbol{z}}_{H}$. In this case, we can eliminate the redundant latent encoding/decoding in computing the regularization loss, as we follow DMD [58] to optimize the distribution loss in the latent state space rather than in the noise space. The gradient of the regularization loss w.r.t. the network parameter $\theta$ in the latent space is:
@@ 4.1 Experimental Settings
Training and Testing Datasets. Prior works [42, 57, 31, 52] employed different training datasets, making unified training standards for Real-ISR difficult to establish. For simplicity, we adopt SeeSR’s setup [52] and train OSEDiff using the LSDIR [26] dataset and the first 10K face images from FFHQ [19]. The degradation pipeline of Real-ESRGAN [45] is used to synthesize LQ-HQ training pairs. We evaluate OSEDiff and compare it with competing methods using the test set provided by StableSR [42], including both synthetic and real-world data. The synthetic data includes 3000 images of size $512\times 512$, whose GT are randomly cropped from DIV2K-Val [2] and degraded using the Real-ESRGAN pipeline [45]. The real-world data include LQ-HQ pairs from RealSR [3] and DRealSR [51], with sizes of $128\times 128$ and $512\times 512$, respectively.
@@
Inference Time (s) 11.50 2.72 4.30 2.80 0.71 0.13 0.11
MACs (G) 79940 24234 65857 29125 5491 2649 2265
# Total Param (M) 1410 1717 2524 1900 119 119 1775
@@
Evaluation Metrics. To provide a comprehensive and holistic assessment on the performance of different methods, we employ a range of full-reference and no-reference metrics. PSNR and SSIM [50] (calculated on the Y channel in YCbCr space) are reference-based fidelity measures, while LPIPS [64], DISTS [12] are reference-based perceptual quality measures. FID [15] evaluates the distance of distributions between GT and restored images. NIQE [62], MANIQA-pipal [55], MUSIQ [22], and CLIPIQA [41] are no-reference image quality measures. We also conduct a user study, which is presented in the Appendix.
@@ 4.2 Comparison with State-of-the-Arts
Complexity Comparisons. We further compare the complexity of competing DM-based Real-ISR models in Table 2, including the number of inference steps, inference time, and trainable parameters. All methods are tested on an A100 GPU with an input image of size $512\times 512$. OSEDiff has the fewest trainable parameters, and the trained LoRA layers can be merged into the original SD to further reduce the computational cost. By using only one forward pass, OSEDiff has significant advantage in inference time over multi-step methods. Specifically, OSEDiff demonstrates a substantial speed advantage, being approximately 105 times faster than StableSR, 39 times faster than SeeSR, and 6 times faster than ResShift. When compared to the single-step method SinSR, OSEDiff not only achieves faster inference but also delivers significantly higher output quality. In terms of complexity, OSEDiff requires the lowest MACs at just 2265G, as it operates with only a single diffusion step. In contrast, methods like StableSR, which require 200 steps, incur substantially higher MACs (e.g., 79940G). Regarding trainable parameters, OSEDiff is highly parameter-efficient, requiring only 8.5M parameters (LoRA layers), compared to models such as SeeSR, which necessitates 749.9M parameters. This highlights the efficiency of OSEDiff during the training process.
@@ Methods
Table 7: Ablation studies on finetuning the VAE encoder and decoder on the RealSR benchmark.
Train VAE Encoder Train VAE Decoder PSNR↑ DISTS↓ LPIPS↓ CLIPIQA↑ MUSIQ↑ NIQE↓
(1) × × 25.27 0.1966 0.2656 0.5303 58.99 6.5496
(2) × 25.30 0.2049 0.2829 0.5604 65.83 6.6291
(3) 25.59 0.2141 0.3017 0.5778 65.92 6.9845
OSEDiff × 25.15 0.2128 0.2921 0.6693 69.09 5.6479
@@ 4.3 Ablation Study
The Finetuning on the VAE Encoder and Decoder. We conducted ablation studies to examine the impact of finetuning the VAE encoder and decoder, as shown Table 7. In the first row, where neither the VAE encoder nor decoder is finetuned, the results show poor perception performance. Comparing with OSEDiff, where only the VAE encoder is finetuned, we observe significant improvements in perceptual quality (e.g., MUSIQ improves from 58.99 to 69.09). This demonstrates that finetuning the VAE encoder is important for removing degradation and enhancing overall performance. When comparing the third row, where both the VAE encoder and decoder are finetuned, with OSEDiff, where only the encoder is trained and the decoder is fixed, we note that OSEDiff also achieves better perceptual quality (CLIPIQ improves from 0.5778 to 0.6693). This indicates that fixing the VAE decoder is important to ensure that the UNet output remains in the original VAE latent space, which helps minimizing the VSD loss more effectively. Thus, finetuning the VAE encoder is important to remove degradation, while fixing the VAE decoder helps maintaining stability in the latent space, leading to better perceptual quality.
@@ A.4 Algorithm of OSEDiff
The pseudo-code of our OSEDiff training algorithm is summarized as Algorithm 1. We follow [49, 58] and use classifier-free guidance (cfg) when calculating $\boldsymbol{z}_{\phi}$.The cfg value is set to 7.5, and the negative prompts we use are: "painting, oil painting, illustration, drawing, art, sketch, oil painting, cartoon, CG Style, 3D render, unreal engine, blurring, dirty, messy, worst quality, low quality, frames, watermark, signature, jpeg artifacts, deformed, lowres, over-smooth."
@@ 14      $\hat{\boldsymbol{x}}_{H}\leftarrow\emph{D}_{\theta}(\hat{\boldsymbol{z}}_{H})$
16      $\nabla_{\theta}\mathcal{L}_{\mathrm{data}}\leftarrow\left[\mathcal{L}_{\mathrm{MSE}}(\hat{\boldsymbol{x}}_{H},\boldsymbol{x}_{H})+\lambda_{1}\mathcal{L}_{\mathrm{LPIPS}}(\hat{\boldsymbol{x}}_{H},\boldsymbol{x}_{H})\right]\frac{\partial\hat{\boldsymbol{x}}_{H}}{\partial\theta}$
      /* Compute regularization objective, following DMD[58] */
@@ 18      Sample $\epsilon$ from $\mathcal{N}(0,\boldsymbol{I})$
19      Sample $t$ from $\{20,\cdots,980\}$
21      $\boldsymbol{z}_{\phi}\leftarrow\operatorname{stopgrad}(F_{\phi}\left(\hat{\boldsymbol{z}}_{t};c_{y}\right))$
22      $\boldsymbol{z}_{\phi^{\prime}}\leftarrow\operatorname{stopgrad}(F_{\phi^{\prime}}\left(\hat{\boldsymbol{z}}_{t};c_{y}\right))$
23      $\omega\leftarrow 1/\mathrm{mean}(\|\boldsymbol{z}_{\phi}-\hat{\boldsymbol{z}}_{H}\|)$
24      $\nabla_{\theta}\mathcal{L}_{\mathrm{reg}}\leftarrow\left[\omega(\boldsymbol{z}_{\phi^{\prime}}-\boldsymbol{z}_{\phi})\right]\frac{\partial\hat{\boldsymbol{z}}_{H}}{\partial\theta}$
      /* Compute regularizer funetuning objective */
25
26      Sample $\epsilon$ from $\mathcal{N}(0,\boldsymbol{I})$
@@ 27      Sample $t$ from $\{1,\cdots,T\}$
28      $\boldsymbol{z}_{t}\leftarrow\alpha_{t}\operatorname{stopgrad}(\hat{\boldsymbol{z}}_{H})+\sigma_{t}\epsilon$
29      $\mathcal{L}_{\mathrm{diff}}\leftarrow\mathcal{L}_{\mathrm{MSE}}(\boldsymbol{\epsilon}_{\phi^{\prime}}(\boldsymbol{z}_{t};t,c_{y}),\epsilon)$